Suppose that a random variable X has a DISCRETE distribution with probability mass function given as: P(x) = ( x / 2485 if x = 1, 2, . . . , 70 0 Otherwise a) Find P(x > 5) b) Find P(2 ≤ x ≤ 68) c) Find P(x = 70.5)


1
Expert's answer
2021-10-01T15:22:32-0400

Solution.

"Р(x)=\begin{Bmatrix}
 \frac{x}{2485}, \text{if }x=1,2,...,70 \\
 0, otherwise
\end{Bmatrix}"

a)

"P(x>5)=1-P(x\leq 5)=1-\frac{1}{2485}-\frac{2}{2485}-\frac{3}{2485}-\frac{4}{2485}-\frac{5}{2485}=1-\frac{15}{2485}=\frac{2470}{2485}=0.994"

b)

"P(2\leq x\leq 68)=P(68)-P(2)=\frac{68}{2485}-\frac{2}{2485}=\frac{66}{2485}=0.027"

c)

70.5>70, so 70.5 in otherwise.

"P(x=70.5)=0"